OpenStudy (anonymous):

Find all rational root of the following function: f(x)= -2x(x-6)(x+1)

OpenStudy (triciaal):

f(x) = 0 -2x = 0 ; x- 6 = 0 ; x + 1 = 0 0, -1, 6

OpenStudy (anonymous):

Can you explain how to find the rational roots?

OpenStudy (triciaal):

0 is not rational

OpenStudy (anonymous):

So would it just be 6 then?

OpenStudy (triciaal):

0 is not rational

OpenStudy (triciaal):

the roots are the values of x that make f(x) = 0 same as where the function/graph crosses the x-axis

OpenStudy (triciaal):

if (x-a) is a factor then there is no remainder when x = a